Status Report

NASA Space Station On-Orbit Status 22 July 2004

By SpaceRef Editor
July 27, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 22 July 2004

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2004) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally except those noted previously or below.
CDR Padalka completed the Russian software upgrade, today loading the no. 2 Central Post computer (KTsP2), Laptop 2 (LT2) and Laptop 3 (LT3) with the new version 7.02.   [Installation of 7.02 on KTsP2 was performed from LT2 using the older 7.01 software.  LT2 was afterwards upgraded itself, followed by LT3.  Both LTs were reconfigured for running Windows.  The task also included checkout of the LT2-to-KTsP2-to-TsVM data linkage, and transfer of BVS (SM computer system) memory images to KTsP1.]

In preparation for the upcoming Orlan EVA-10 (8/3) FE/SO Fincke retrieved a spare portable air repress bottle (BNP) from the “divan” of the Soyuz orbital module and installed it in the DC1 docking compartment’s instrumentation “corridor”.   [This additional BNP is used as standard protection against the contingency of the repress valve to the corridor stuck Closed during the post-ingress repressurization of the DC1.]

Later, Fincke also installed a BNP air tank in the Service Module (SM)’s work compartment (RO) repress lines.

Previous Reports

ISS On-orbit Status [HQ]
ISS Status [JSC]
Shuttle Processing [KSC]

The crew began unstowing and gathering the consumable/replaceable ORU elements for the Orlan-M suits to be used on 8/3.  They also set up Laptop 9 with the Orlan/EVA training software.   [As for EVA-9, the CDR will be wearing Orlan #25 (red markings) with BRTA radio telemetry unit #13 installed, while the FE’s suit will be #26 (blue markings) with BRTA #18.  Orlan ORUs are LiOH canisters (LP-9), primary & backup oxygen tanks (BK-3), moisture collectors, feedwater filters (FOR), CO2 measuring unit (IK) filter, filtration & separation units (BOS), and 825M1 “akkumulyatoriy” (storage batteries).]

Padalka and Fincke took turns in performing the mandatory Russian pre-EVA MedOps procedure MO-6 (hand-cycle ergometry), each providing pomosh’ (assistance) to the other during the individual’s 30-min. test run.   [Because previous cosmonauts have shown noticeable decrease in arm muscle tone, TsUP/IBMP (MCC-Moscow/Institute of Biomedical Problems) physical fitness experts have made the handgrip/arm tolerance test analysis (hand ergometry) a standard pre-Orlan EVA requirement.  MO-6 today required the CDR, later the FE, to don the ECG (electrocardiogram) biomed harness, attach three skin electrodes and plug the harness into the PKO medical exam panel on the cycle ergometer.  The exercise itself started after 10 seconds of complete rest, by manually rotating the cycle’s pedals, set at 150 W, backwards until “complete exhaustion”.  It was supported by tagup with ground specialists.]

At ~6:15am EDT, the CDR held his weekly IMS tagup with ground specialists, discussing open issues concerning identification of equipment and storage locations for updating the IMS database, via S-band.  Today’s exchange included a calldown of numbers of trash bags (KBO) stowed in Progress 14P.

Mike performed the regular once-a-week maintenance reboot on the operational PCS laptops and also restarted the OCA comm router laptop (every two weeks).

The FE also collected the periodic CO2 partial pressure reading in the SM and Lab using the U.S. CDMK (CO2 monitor kit), for calldown to MCC-Houston (along with battery status) for use in trending analyses.

Later in the day Fincke terminated the maintenance recharge of the #2045 & #2046 EMU batteries in the Airlock (A/L)’s battery stowage assembly (BSA), then stored them in the A/L after resetting their 50-day clock.

At ~1:05pm EDT, Mike started the DOUG (Dynamic Operational Ubiquitous Graphics) software at the Lab RWS (Robotics Work Station), and the crew subsequently operated the Canadian SSRMS (Space Station Remote Manipulator System), moving it in the position best suited for video-observation of the 8/3 spacewalk.

Gennady completed the daily routine inspection of the SM’s SOZh life support system and prepared the daily IMS (inventory management system) “delta” file update.

Loading of US trash on 14P has been completed, including 15 food containers.

At 9:00am, the crew On DO3, at 4:30am EDT, the crew downlinked a TV message of greetings to the Third International Children’s Games, “Children of Asia”, in Yakutsk in the Republic of Sakha (Yakutia).  [These third international sports games for kids are held by Government of Sakha’s Federal Agency for Physical Education, Sports and Tourism along with the Russian Olympic Committee on July 23 through 30, 2004.]

The crew completed their daily 2.5-hr. physical exercise program on TVIS, RED exerciser and VELO cycle with load trainer.

Later tonight Fincke is scheduled to set up the amateur radio gear and conduct a voice check of the Kenwood equipment during a ground station pass.

Before sleep time, Mike will be briefed by EarthKAM (Earth Knowledge Acquired by Middle School Students) ground personnel on the results of the recent highly successful EK run.

For his next “Saturday Science” program, the Science Officer has selected another session with the ISSI (In-Space Soldering Investigation) experiment.   [The ISSI PI has reported that previous ISSI photography is exhibiting unexpected and unusual features of the soldered coupons.]

On 7/19 the U.S. starboard thermal radiator (STR) MDM computer failed “hard”, as a subsequent investigation confirmed.  The MDM was powered off, which also means loss of redundancy for the RBVM (Radiator Beam Valve Modules) fluid line heaters.   [These heaters thermally condition the flexhose, manifold lines and QDs (quick disconnects) running along the length of the STR beam.  The STR is not yet charged with ammonia (NH3) and will not be needed before Flight 12A.1.  Thus the MDM failure is of minor impact.  If it is not recoverable, it will have to be replaced by EVA.  A spare unit is available on board.]

TsUP/Moscow reports that two EMU water pumps have formally been included in the Progress 15P manifest.

Moscow and Houston have jointly worked out an oxygen (O2) repressurization plan for the ISS, which uses the 14 kg O2 still stored on Progress 14P.  For the time being, Moscow considers the O2 an emergency supply for contingencies until shortly before 14P undocking.  Then, it will be used to repress the station, raising its O2 pressure to the upper Flight Rule limit.  The Elektron will continue to run at its lowest setting (16 amp).  [The repress amounts to a delta O2 pressure increase of 20 mmHg (Torr), increasing O2 pressure to 180 mmHg or ~23.9% ppO2.  The gas repress will probably be conducted in two refreshes of ~9 mmHg each since there is no prior experience with a 20 mmHg push.]

On Tuesday (7/20), the Elektron shut down temporarily, probably because of another bout with air bubbles.  It was restarted yesterday and is working nominally.
Mike Fincke’s ultrasound investigation of the Lap window for leaks yesterday has not evidenced any unusual leak sounds beyond the ambient sound level.

Correction of yesterday’s status:  The “Eagle” of Apollo 11 touched down on Mare Tranquillitatis, of course (not on M. Serenitatis).  Another “shame on me”.

Major upcoming events:

  • Progress 14P undocking — 7/30 (2:06am EDT);
  • Orlan EVA-10 from DC-1 docking compartment — 8/3 (hatch open: 2:50am EDT);
  • Progress 15P launch — 8/11 (1:01am EDT);
  • Progress 15P docking — 8/14 (2:05am EDT).

Expedition 9 Flight Crew Plans can be found at

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

Today’s CEO (crew earth observation) photo targets, limited in the current XPOP attitude by flight rule constraints on the use of the science window, which is available for only ~1/4 of each orbit when not facing forward (in “ram”), were Betsiboka River Delta, Madagascar (nadir pass.  Prior ISS/CEO imagery has largely answered the science objective.  But detailed imagery is so good that investigators can now examine the more detailed question of coastline positions at high and low tide, since the estuary is so shallow), Amman, Jordan (looking a touch left), Athens, Greece (nadir pass), and Internal waves, Pacific coast, Central America (looking left toward the glint point.  The most interesting zone closest to shore is predicted to be cloud free).

CEO images can be viewed at these websites:

See also the website “Space Station Challenge” at:

To view the latest photos taken by the expedition 9 crew visit:

U.S. & Russian Segment Status  (as of today, 1:17pm EDT)

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is On (16A, =lowest setting).  Vozdukh CO2 scrubber is On.  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is operating.  SM Gas Analyzer has been calibrated and is used for ppO2 & ppCO2 monitoring.  MCA (major constituents analyzer) is in Life Extending Mode (LEM).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On, SKV-2 is Off (SM panel mods completed 4/8; SKV-2 activation failed 4/20; is still considered failed).  SFOG slot #2 fan suspect (not usable).
  • SM Working Compartment:  Pressure (mmHg) — 753; temperature (deg C) — 25.9; ppO2 (mmHg) — 161.0; ppCO2 (mmHg) — 4.0.
  • SM Transfer Compartment:  Pressure (mmHg) — 751; temperature (deg C) — 20.2.
  • FGB Cabin:  Pressure (mmHg) — 752; temperature (deg C) — 22.0.
  • Node:  Pressure (mmHg) — 743.9; temperature (deg C) — 23.5 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 745.3; temperature (deg C) — 24.2; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 745.5; temperature (deg C) — 25.5; shell heater temp (deg C) — n/a, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.

(n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in Autotrack (solar-tracking, “sun slicer”, i.e., drag reduction-biased by 47 deg).
  • SM batteries:  All batteries (8) are on line in “Partial Charge” mode.  
  • FGB batteries:  Battery #4 is off line; battery 5 is on line in “Cycle” mode; all other batteries (4) are on line in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is backup, and C&C-3 is in standby.
  • GNC-2 MDM is prime; GNC-1 is backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-1 is On (primary), EXT-2 is Off (backup).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.
  • FGB MDM-1 is Off (failed, 11/21/03); MDM-2 is Operational.

Propulsion System:

  • Total propellant load available: 3864 kg (8519 lb) as of 7/15/04;  [SM(552) + FGB(3312) + Progress M(0)].  (Capabilities: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed, since 6/6/02; CMG-2’s RPC-17 failed 4/21/04; was replaced 6/30/04).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • XPOP (x-axis perpendicular to orbit plane = “sun-fixed” [yaw: 0.5 deg, pitch: -9.0 deg., roll: 0 deg]), with CMG TA (thruster assist), until 7/26, then LVLH XVV.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally (may require a mask).
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, operational on redundant string, off on prime.
  • MBS: KA (keep alive) power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Location NOW

Full Size/Update
Real Time ISS TrackerMore Links

ISS Orbit  (as of this morning, 6:30am EDT [= epoch]):

  • Mean altitude — 358.9 km
  • Apogee height — 363.1 km
  • Perigee height — 354.8 km
  • Period — 91.7 min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.0006175
  • Solar Beta Angle — 31.5 deg
  • Orbits per 24-hr. day — 15.70
  • Mean altitude loss in last 24 hours — 108 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 32385

ISS Altitude History

Apogee height Mean AltitudePerigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at

SpaceRef staff editor.